Alex has $2.85 in coins in his pocket. He has 3 times as many quarters as dimes and half as many dimes as nickels. How many nickels does he have in his pocket?
Found 2 solutions by phoihe001, y6n3o9p4:
Answer by phoihe001(34)
Answer by y6n3o9p4(20)

The amount of quarters must be a multiple of 3 because otherwise that would mean the amount of dimes is a decimal, so try plugging in the numbers with the amount of quarters being a multiple of 3 until the total value of the coins is $2.85.
q = $0.25, d = $0.10, n = $0.05
3q+d+2n = $0.95 (Not enough)
6q+2d+4n = $1.90 (Not enough)
9q+3d+6n = $2.85 (Correct)
Alex has 6 nickels in his pocket.